OpenStudy (anonymous):

The graph shows the solution to the initial value problem y'(t)=mt, y(t0)= -4 find: m= t0= y(t)=

OpenStudy (amistre64):

x^2-4? or t^2 - 4?

OpenStudy (anonymous):

t^2-4

OpenStudy (turingtest):

why do you say "presumably" y=t^2-4 ???

OpenStudy (turingtest):

there seems to be some missing info here, but you can think of this as\[y'(t)=mt\implies y(t)=\int mtdt=\frac m2t^2+y(t_0)\]which we can use to find m and \(y(t_0)\) if we know the graph